# Periodic classical trajectories and quantum scars in many-spin systems

### v3

@@ I Introduction
The role and the properties of chaos in the process of dynamic thermalization of many-body systems remain a subject of intense research that keeps resisting generic conjectures[1, 2, 3]. The open issues concern the transition from classical to quantum thermalization[4, 5], as well as the thermalization in purely classical systems as such[6, 7]. Central to those issues is the lack of a clear border between systems exhibiting and not exhibiting dynamic thermalization to the Gibbs equilibrium. It is often the case that a seemingly exotic non-thermalising setting, in fact, sets a warning post for a much broader class of physical situations, where the thermalization is absent or poor. In the latter case, the system may be thermalisable in principle but only on a rather long timescale[8, 9, 10, 11, 12, 13, 14, 15, 16]. The examples of nonthermalising/poorly thermalising behavior are: integrable systems [17, 18, 19, 20], glassy systems [21, 22], systems possibly exhibiting many-body localisation[23, 24, 25, 26, 27], periodically driven systems[28, 29, 30, 31, 32, 33, 34, 35, 36], and, finally, coming to the topic of this article, systems with quantum many-body scars [37, 38, 39].
@@ II.2 Quantum scars
A recent addition to the zoo of poorly thermalising settings is the case of the so-called “quantum many-body scars” associated with anomalously slow relaxation of quantum lattice systems that was first observed experimentally in Ref.[38] and then interpreted theoretically in Ref.[39] in terms of the “PXP” quantum model for a spin-1/2 chain. It was found that the PXP model possesses atypical eigenstates inconsistent with eigenstate thermalization[53, 54], and, at the same time, these atypical eigenstates are prominently present in the initial nonequilibrium state of the experimentally observed system of Ref.[38]. Eigenstate thermalization hypothesis (ETH) stipulates that the quantum expectation values of a physical observable in a normally thermalising system remain almost the same for all energy eigenstates within a given energy window. The atypical eigenstates of the PXP model named “quantum scars” stood out in this respect. A similar kind of ETH violation was, actually, reported earlier in Ref. [37] without reference to quantum scars. The above findings were followed a large body of research [106, 107, 108, 109, 110, 111, 112, 113, 114].
@@
Let us recall here that the concept of quantum scars as such was initially introduced by Heller[115] in the eighties for one particle moving in a chaotic two-dimensional billiard: it referred to atypical quantum eigenstates corresponding to classical periodic trajectories with short periods. When a classical periodic trajectory is sufficiently short, then the quasiclassical limit of the quantum dynamics implies the onset of quantum interference corrections associated with the trajectory closing onto itself before the chaotic Lyapunov instability suppresses such interference, and these corrections, in turn, amplify the probability of the corresponding eigenstate to remain localized around the classical trajectory as opposed to being evenly spread over the entire system which the typical eigenstates do. Numerous examples of low-dimensional quantum scars have been investigated since then – see, e.g., [116, 117]. Yet, a missing link between the original low-dimensional quantum scars and the recently discovered quantum many-body scars was that the latter were not connected to periodic classical trajectories. In other words, the quantum many-body scars as discovered were rather “distant relatives” to the quantum scars known from the physics of billiards. This issue of missing periodic trajectories was addressed in [118, 119], although not in the classical limit but rather in the parameter space of quantum matrix product states.
@@ III Hamiltonian and initial conditions
[…]
@@
Here delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over all possible values of spin projections, quantum or classical, which implies that $\langle S_{i}^{\alpha}\rangle=0$ and $\langle S_{i}^{\alpha}S_{i+1}^{\alpha}\rangle=0$. Further noting the equivalence between the canonical and the microcanonical ensembles in the limit $L\to\infty$, we conclude that an initial spin state with zero energy belongs to the energy shell of maximum entropy (maximum density of states quantum mechanically).
@@ IV.1 Equations of motion
[…]
@@ IV.4 Lyapunov instabilities: general considerations
Generically, different wave vectors $\mathbf{q}_{k}$ are supposed to correspond to different values of $\lambda_{k}$ with the exception of symmetry-related or accidental degeneracies. A symmetry-related degeneracy occurs, in particular, when a lattice, such as the one we consider, has inversion symmetry, and, as a result, pairs of wave vectors ±$\pm\mathbf{q}_{k}$ correspond to the same $\lambda_{k}$. In this case, a generic Lyapunov exponent is at least double-degenerate, the possible non-degenerate cases correspond to $\mathbf{q}_{k}=0$ and $\mathbf{q}_{k}=\pi/L$.
From now on, we focus on our one-dimensional lattice, where symmetry representations are characterised by wave numbers $q$ rather than wave vectors, and so are the Lyapunov instabilities. The above discussion implies that, when the periodic trajectory $\mathcal{S}_{\text{p}}(t)$ is Lyapunov unstable, the instability corresponding to the largest Lyapunov exponent $\lambda_{\text{p}}$ generically develops out of the initially random and hence not translationally invariant deviation $\delta\mathcal{S}(0)$ as a standing wave in real space characterised by a pair of wave numbers ±$\pm q_{\text{p}}$.
@@ IV.5.1 Overview
Figure 5: Spatial Fourier spectra $f_{\lambda_{\text{p}}}(q)$ of the Lyapunov vectors defined by Eq. (12). The chain lengths $L$ and the coupling constants $J$ are indicated in plot legends. The allowed values of wavenumbers are $q=\frac{2\pi}{L}k$, where $k$ is an integer in the interval $[-L/2,L/2]$, $t_{0}=750$.
@@
of the Lyapunov vectors $\delta{\cal S}(t)$ corresponding to $\lambda_{\text{p}}$ with time $t$ being large enough for the computed $\lambda_{\text{p}}$ to reach the stationary value. Here $\delta\mathbf{S}_{m}\equiv(\delta S_{m}^{x},\delta S_{m}^{y},\delta S_{m}^{z})$ are the components of the high-dimensional vector $\delta{\cal S}$ entering Eq.(11); the Fourier amplitude inside the absolute-value bars |||...|| … | is a three-dimensional vector with $x$, $y$ and $z$ complex-valued projections, $|...|^{2}$ is the sum of the squares of the moduli of those projections. The Fourier spectra in Fig. 5 are computed for two chain lengths $L=18$ and $L=20$ and for two interaction constants $J$ representing the libration and the rotation regimes for each length. In all four cases, one can see two dominant Fourier peaks corresponding to ±$\pm q_{\text{p}}$.
@@
We conjecture that the oscillatory dependence of $\lambda_{\text{p}}$ on $L$ in Fig.4, originates from the fact that, for a chain of a given length $L$, the allowed values of the wave number $q=\frac{2\pi}{L}n$, where $n=0,1,...,L-1$, cannot exactly match a certain value $q_{0}$ that is optimal for maximising the Lyapunov exponent $\lambda_{\text{p}}$ in a chain of infinite length. Let us denote that maximal value as $\lambda_{\text{p,max}}$. As the length of the chain increases, the allowed values of $q$ come increasingly closer to $q_{0}$, and, as a result, the oscillatory part of $\lambda_{\text{p}}(L)$ becomes gradually suppressed. We are dealing with a supposedly small difference between $q_{0}$ and the nearest allowed value $q_{\text{p}}=\frac{2\pi}{L}\ \text{round}\left(\frac{q_{0}L}{2\pi}\right)$, where function “$\text{round}(...)$” rounds its argument to the nearest integer value. It is thus reasonable to parameterise the departure from $\lambda_{\text{p,max}}$ by a quadratic dependence on $(q_{0}-q_{\text{p}})$:
@@
$\lambda_{\text{p}}(J)$ $J\to J^{*}\pm 0]{}$ $\lambda_{A};$ (14)
[…]
@@
Figure 8: Deviation of a slightly perturbed ensemble-averaged dynamics of the $z$-projection of one spin, $\bar{S}^{z}_{1}(t)$, from the maximum value 1 corresponding to a perfectly periodic motion. The inset displays $\bar{S}^{z}_{1}(t)$ (blue solid line). This plot was computed for a spin chain of length $L=100$ with interaction parameter ${J=1.76}$. The averaging was performed over the ensemble of $N=1000$ realisations of $S^{z}_{1}(t)$ with initial conditions ${\mathbf{S}_{n}(0)=\left(\delta S_{n0}^{x},\delta S_{n0}^{y},\sqrt{1-\delta{S_{n0}^{x}}^{\!\!{}^{2}}-\delta{S_{n0}^{y}}^{\!\!{}^{2}}}\right)}$, where $\delta S_{n0}^{x}$ and $\delta S_{n0}^{y}$ were randomly sampled from within the circle $\sqrt{\delta{S_{n0}^{x}}^{\!\!{}^{2}}+\delta{S_{n0}^{y}}^{\!\!{}^{2}}}\leq 10^{-4}$. The main panel is a semilog plot of the points $1-\bar{S}^{z}_{1}(t_{m})$, where times $t_{m}$ correspond to the maxima of $\bar{S}^{z}_{1}(t_{m})$ plotted in the inset. These points are represented by the red dots in both the main frame and the inset. The highlighted time interval marks the Lyapunov regime of the exponential growth of $1-\bar{S}^{z}_{1}(t_{\text{m}})$. The green solid line is the exponential fit $\exp(2\lambda_{p}t_{\text{m}})$, where $\lambda_{p}=0.303$ is the value directly computed in the many-spin phase space. [The initial scatter of the points in the main panel is due to the mismatch between the discretised time points and the true maxima of $\bar{S}^{z}_{1}(t)$.]
@@ V Classical spins: transient nearly quasiperiodic regime
While computing the Lyapunov exponents for periodic trajectories, we discovered that, as the Lyapunov instability develops, the deviations from the reference periodic trajectory become large and no longer describable by Lyapunov exponents, and yet the dynamics does not immediately become ergodic — rather it often enters a nearly quasiperiodic regime illustrated in Fig. 9. The lifetime of this regime exhibits strong variations: it may be relatively short — a few time units (one unit is $11/h=$ 1), or fairly long — a few hundred time units — as is the case in Fig. 9(b), or anomalously long — meaning $\sim 10^{6}$ time units or more; one such example is shown in Figs. 9(a) and 10. Later in this section we identify the anomalously long-lived quasiperiodic regime with the phenomenon of Arnold diffusion. In our simulations we came across only five cases of this regime, which are marked as such in Figs.4(b,c).
@@ V.1 Frequency spectra
[…]
@@ V.4 Time evolution of Fourier modes with different wave numbers $q$
We further note that the exponential onset of each $q=q_{p}$ mode in Fig. 12 is accompanied by the subsequent onset of the modes growing as $F_{q}(t)\propto\exp({2n\lambda_{\text{p}}t})$, where $n$ is a natural number. These modes are explicitly indicated by the respective exponential fits in Fig. 12(c). The straight lines representing these exponential fits on a semilog plot tend to intersect roughly at the same point located on the red line representing the mode $F_{0}(t)$. This phenomenology can be explained as follows. The growth of the leading $q=q_{p}$ mode induces a modulation of the chain background, which, on the one hand, constitutes a dominant perturbation to the original translationally invariant periodic trajectory but, on the other hand, still remains very small as long as $F_{q_{p}}(t)\ll F_{0}(t)$. As a result, the spatial Fourier harmonics with $q=nq_{p}$ become coupled to the $q=q_{p}$ mode in the $n$th order of the perturbation theory, and their amplitudes emerge at later times out of statistical noise exhibiting the Lyapunov growth with the rate $n\lambda_{p}$, [$2n\lambda_{p}$ for $F_{q}(t)$]. This phenomenology, however, does not signify independent modes – rather they are perturbation-induced admixtures to the original unstable $q=q_{p}$ mode. That perturbation can be described in terms of backfolding of the chain’s Brillouin zone, which is to be discussed further below in subsection V.6 in the context of Mechanism B.
@@ V.6 Mechanism B and the backfolding of the Brillouin zone
In order to classify these possible additional instabilities, we note that, since the quasiperiodic motion involves a $q_{\text{p}}\neq 0$ mode, the setting of the respective Lyapunov problem is no longer translationally invariant with respect to the spin chain in the sense that the instantaneous spin pattern does not repeat under the translation $\mathbf{S}_{m}\to\mathbf{S}_{m+1}$ (periodically closed). Moreover, the wave number $q_{\text{p}}=2\pi n/L$, where $n$ is an integer, corresponds to the spatial wavelength $\Lambda_{\text{p}}\equiv{2\pi\over q_{\text{p}}}=L/n$, which is, in general, not an integer number and thus cannot be a period with which a spin pattern repeats itself. That period is rather $m\Lambda_{\text{p}}$, where $m$ is the minimal natural number that makes $m\Lambda_{\text{p}}=mL/n$ an integer. This implies that $u\equiv n/m$ should be the largest common divider of $L$ and $n$. The spatial period of the $q_{\text{p}}$-mode is thus $L/u$, while the corresponding wave number determining the size of the reduced Brillouin zone of the two-mode dynamics is $Q_{u}=2\pi u/L$. The wave numbers from the original Brillouin zone of the chain $[-\pi,\pi]$ should now be backfolded into the new reduced Brillouin zone $[-Q_{u}/2,Q_{u}/2]$. The backfolded modes then form $L/u$ “bands” of Lyapunov exponents.
@@
When $u>1$, the reduced Brillouin zone has $u$ different $q$ points – each representing a group of the backfolded $q$ values, in which case the modes belonging to the group containing the wave vectors ±$\pm q_{\text{p}}$ become perturbatively coupled to the $q_{\text{p}}$-mode, while the rest of the modes remain uninvolved.

### v4

# Classical periodic trajectories and quantum scars in many-spin systems

@@ I Introduction
The role and the properties of chaos in the process of dynamic thermalization of many-body systems remain a subject of intense research that keeps resisting generic conjectures[1, 2, 3]. The open issues concern the transition from classical to quantum thermalization[4, 5], as well as the thermalization in purely classical systems as such[6, 7]. Central to those issues is the lack of a clear border between systems exhibiting and not exhibiting dynamic thermalization to the Gibbs equilibrium. It is often the case that a seemingly exotic non-thermalizing setting, in fact, sets a warning post for a much broader class of physical situations, where the thermalization is absent or poor. In the latter case, the system may be thermalizable in principle but only on a rather long timescale[8, 9, 10, 11, 12, 13, 14, 15, 16]. The examples of nonthermalizing/poorly thermalizing behavior are: integrable systems [17, 18, 19, 20], glassy systems [21, 22], systems possibly exhibiting many-body localisation[23, 24, 25, 26, 27], periodically driven systems[28, 29, 30, 31, 32, 33, 34, 35, 36], and, finally, coming to the topic of this article, systems with quantum many-body scars [37, 38, 39].
@@ II.2 Quantum scars
A recent addition to the zoo of poorly thermalizing settings is the case of the so-called “quantum many-body scars” associated with anomalously slow relaxation of quantum lattice systems that was first observed experimentally in Ref.[38] and then interpreted theoretically in Ref.[39] in terms of the “PXP” quantum model for a spin-1/2 chain. It was found that the PXP model possesses atypical eigenstates inconsistent with eigenstate thermalization[53, 54], and, at the same time, these atypical eigenstates are prominently present in the initial nonequilibrium state of the experimentally observed system of Ref.[38]. Eigenstate thermalization hypothesis (ETH) stipulates that the quantum expectation values of a physical observable in a normally thermalizing system remain almost the same for all energy eigenstates within a given energy window. The atypical eigenstates of the PXP model named “quantum scars” stood out in this respect. A similar kind of ETH violation was, actually, reported earlier in Ref. [37] without reference to quantum scars. The above findings were followed a large body of research [106, 107, 108, 109, 110, 111, 112, 113, 114].
@@
Let us recall here that the concept of quantum scars as such was initially introduced by Heller[115] in the eighties for one particle moving in a chaotic two-dimensional billiard: it referred to atypical quantum eigenstates corresponding to classical periodic trajectories with short periods. When a classical periodic trajectory is sufficiently short, then the quasiclassical limit of the quantum dynamics implies the onset of quantum interference corrections associated with the trajectory closing onto itself before the chaotic Lyapunov instability suppresses such interference[116], and these corrections, in turn, amplify the probability of the corresponding eigenstate to remain localized around the classical trajectory as opposed to being evenly spread over the entire system which the typical eigenstates do. Numerous examples of low-dimensional quantum scars have been investigated since then – see, e.g., [117, 118]. Yet, a missing link between the original low-dimensional quantum scars and the recently discovered quantum many-body scars was that the latter were not connected to periodic classical trajectories. In other words, the quantum many-body scars as discovered were rather “distant relatives” to the quantum scars known from the physics of billiards. This issue of missing periodic trajectories was addressed in [119, 120], although not in the classical limit but rather in the parameter space of quantum matrix product states.
@@ III Hamiltonian and initial conditions
In all concrete calculations throughout the paper, we fix $h=1$, while varying $J$. This implies that the unit of time in all plots is $1/h$.
@@
Here $\langle$$\rangle$ denotes the average over all possible values of spin projections, quantum or classical, which implies that $\langle S_{i}^{\alpha}\rangle=0$ and $\langle S_{i}^{\alpha}S_{i+1}^{\alpha}\rangle=0$. Further noting the equivalence between the canonical and the microcanonical ensembles in the limit $L\to\infty$, we conclude that an initial spin state with zero energy belongs to the energy shell of maximum entropy (maximum density of states quantum mechanically).
@@ IV.1 Equations of motion
where $\{$...,...\} are the Poisson brackets, with $\{S^{\alpha}_{i},S^{\beta}_{j}\}=\delta_{ij}\epsilon_{\alpha\beta\gamma}S^{\gamma}_{i}$. Here $\delta_{ij}$ is the Kronecker tensor, and $\epsilon_{\alpha\beta\gamma}$ is the antisymmetric Levi-Civita tensor. The system of equations (5) can be rewritten as
@@ IV.4 Lyapunov instabilities: general considerations
Generically, different wave vectors $\mathbf{q}_{k}$ are supposed to correspond to different values of $\lambda_{k}$ with the exception of symmetry-related or accidental degeneracies. A symmetry-related degeneracy occurs, in particular, when a lattice, such as the one we consider, has inversion symmetry, and, as a result, pairs of wave vectors $\pm\mathbf{q}_{k}$ correspond to the same $\lambda_{k}$. In this case, a generic Lyapunov exponent is at least double-degenerate, the possible non-degenerate cases correspond to $\mathbf{q}_{k}=0$ and $\mathbf{q}_{k}=\pi/L$.
From now on, we focus on our one-dimensional lattice, where symmetry representations are characterised by wave numbers $q$ rather than wave vectors, and so are the Lyapunov instabilities. The above discussion implies that, when the periodic trajectory $\mathcal{S}_{\text{p}}(t)$ is Lyapunov unstable, the instability corresponding to the largest Lyapunov exponent $\lambda_{\text{p}}$ generically develops out of the initially random and hence not translationally invariant deviation $\delta\mathcal{S}(0)$ as a standing wave in real space characterised by a pair of wave numbers $\pm q_{\text{p}}$.
@@ IV.5.1 Overview
Figure 5: Spatial Fourier spectra $f_{\lambda_{\text{p}}}(q,t=t_{0})$ of the Lyapunov vectors defined by Eq. (12). The chain lengths $L$ and the coupling constants $J$ are indicated in plot legends. The allowed values of wavenumbers are $q=\frac{2\pi}{L}k$, where $k$ is an integer in the interval [L/2,L/2][-L/2,L/2], $t_{0}=750$.
@@
of the Lyapunov vectors $\delta{\cal S}(t)$ corresponding to $\lambda_{\text{p}}$ with time $t$ being large enough for the computed $\lambda_{\text{p}}$ to reach the stationary value. Here $\delta\mathbf{S}_{m}\equiv(\delta S_{m}^{x},\delta S_{m}^{y},\delta S_{m}^{z})$ are the components of the high-dimensional vector $\delta{\cal S}$ entering Eq.(11); the Fourier amplitude inside the absolute-value bars |||...| is a three-dimensional vector with $x$, $y$ and $z$ complex-valued projections, $|...|^{2}$ is the sum of the squares of the moduli of those projections. The Fourier spectra in Fig. 5 are computed for two chain lengths $L=18$ and $L=20$ and for two interaction constants $J$ representing the libration and the rotation regimes for each length. In all four cases, one can see two dominant Fourier peaks corresponding to $\pm q_{\text{p}}$.
@@
[…]
@@
$\lambda_{\text{p}}(J)$ $\xrightarrow[J\to J^{*}\pm 0$]{} $\lambda_{A};$ (14)
$\lambda^{\prime}_{\text{p}}(J)$ $\xrightarrow[J\to J^{*}\pm 0$]{} $\mp\infty.$ (15)
@@
Figure 8: Deviation of a slightly perturbed ensemble-averaged dynamics of the $z$-projection of one spin, $\bar{S}^{z}_{1}(t)$, from the maximum value 1 corresponding to a perfectly periodic motion. The inset displays $\bar{S}^{z}_{1}(t)$ (blue solid line). This plot was computed for a spin chain of length $L=100$ with interaction parameter J=1.76{J=1.76}. The averaging was performed over the ensemble of $N=1000$ realisations of $S^{z}_{1}(t)$ with initial conditions $\mathbf{S}_{n}(0)=\left(\delta S_{n0}^{x},\delta S_{n0}^{y},\sqrt{1-\delta{S_{n0}^{x}}^{\!\!{}^{2}}-\delta{S_{n0}^{y}}^{\!\!{}^{2}}}\right)$}, where $\delta S_{n0}^{x}$ and $\delta S_{n0}^{y}$ were randomly sampled from within the circle $\sqrt{\delta{S_{n0}^{x}}^{\!\!{}^{2}}+\delta{S_{n0}^{y}}^{\!\!{}^{2}}}\leq 10^{-4}$. The main panel is a semilog plot of the points $1-\bar{S}^{z}_{1}(t_{m})$, where times $t_{m}$ correspond to the maxima of $\bar{S}^{z}_{1}(t_{m})$ plotted in the inset. These points are represented by the red dots in both the main frame and the inset. The highlighted time interval marks the Lyapunov regime of the exponential growth of $1-\bar{S}^{z}_{1}(t_{\text{m}})$. The green solid line is the exponential fit $\exp(2\lambda_{p}t_{\text{m}})$, where $\lambda_{p}=0.303$ is the value directly computed in the many-spin phase space. [The initial scatter of the points in the main panel is due to the mismatch between the discretised time points and the true maxima of $\bar{S}^{z}_{1}(t)$.]
@@ V Classical spins: transient nearly quasiperiodic regime
While computing the Lyapunov exponents for periodic trajectories, we discovered that, as the Lyapunov instability develops, the deviations from the reference periodic trajectory become large and no longer describable by Lyapunov exponents, and yet the dynamics does not immediately become ergodic — rather it often enters a nearly quasiperiodic regime illustrated in Fig. 9. The lifetime of this regime exhibits strong variations: it may be relatively short — a few time units (one unit is $1/h=1$), or fairly long — a few hundred time units — as is the case in Fig. 9(b), or anomalously long — meaning $\sim 10^{6}$ time units or more; one such example is shown in Figs. 9(a) and 10. Later in this section we identify the anomalously long-lived quasiperiodic regime with the phenomenon of Arnold diffusion. In our simulations we came across only five cases of this regime, which are marked as such in Figs.4(b,c).
@@ V.1 Frequency spectra
${\cal F}_{1}^{x}(\omega)\equiv\left|\int_{-\infty}^{\infty}S_{1}^{x}(t)e^{-i\omega t}dt\right|^{2}$ (18)
@@ V.4 Time evolution of Fourier modes with different wave numbers $q$
We further note that the exponential onset of each $q=q_{p}$ mode in Fig. 12 is accompanied by the subsequent onset of the modes growing as $F_{q}(t)\propto\exp({2n\lambda_{\text{p}}t})$, where $n$ is a natural number. These modes are explicitly indicated by the respective exponential fits in Fig. 12(c). The straight lines representing these exponential fits on a semilog plot tend to intersect roughly at the same point located on the red line representing the mode $F_{0}(t)$. This phenomenology can be explained as follows. The growth of the leading $q=q_{p}$ mode induces a modulation of the chain background, which, on the one hand, constitutes a dominant perturbation to the original translationally invariant periodic trajectory but, on the other hand, still remains very small as long as $F_{q_{p}}(t)\ll F_{0}(t)$. As a result, the spatial Fourier harmonics with $q=nq_{p}$ become coupled to the $q=q_{p}$ mode in the $n$th order of the perturbation theory, and their amplitudes emerge at later times out of statistical noise exhibiting the Lyapunov growth with the rate $n\lambda_{p}$, $2n\lambda_{p}$ for $F_{q}(t)$]. This phenomenology, however, does not signify independent modes – rather they are perturbation-induced admixtures to the original unstable $q=q_{p}$ mode. That perturbation can be described in terms of backfolding of the chain’s Brillouin zone, which is to be discussed further below in subsection V.6 in the context of Mechanism B.
@@ V.6 Mechanism B and the backfolding of the Brillouin zone
In order to classify these possible additional instabilities, we note that, since the quasiperiodic motion involves a $q_{\text{p}}\neq 0$ mode, the setting of the respective Lyapunov problem is no longer translationally invariant with respect to the spin chain in the sense that the instantaneous spin pattern does not repeat under the translation $\mathbf{S}_{m}\to\mathbf{S}_{m+1}$ (periodically closed). Moreover, the wave number $q_{\text{p}}=2\pi n/L$, where $n$ is an integer, corresponds to the spatial wavelength $\Lambda_{\text{p}}\equiv{2\pi\over q_{\text{p}}}=L/n$, which is, in general, not an integer number and thus cannot be a period with which a spin pattern repeats itself. That period is rather $m\Lambda_{\text{p}}$, where $m$ is the minimal natural number that makes $m\Lambda_{\text{p}}=mL/n$ an integer. This implies that $u\equiv n/m$ should be the largest common divider of $L$ and $n$. The spatial period of the $q_{\text{p}}$-mode is thus $L/u$, while the corresponding wave number determining the size of the reduced Brillouin zone of the two-mode dynamics is $Q_{u}=2\pi u/L$. The wave numbers from the original Brillouin zone of the chain $\pi,\pi]$ should now be backfolded into the new reduced Brillouin zone $Q_{u}/2,Q_{u}/2]$. The backfolded modes then form $L/u$ “bands” of Lyapunov exponents.
@@
When $u>1$, the reduced Brillouin zone has $u$ different $q$ points – each representing a group of the backfolded $q$ values, in which case the modes belonging to the group containing the wave vectors $\pm q_{\text{p}}$ become perturbatively coupled to the $q_{\text{p}}$-mode, while the rest of the modes remain uninvolved.